OpenStudy (anonymous):

f'(x) given f(x)=(2-2x^2)/(2+2x^2)

OpenStudy (anonymous):

pulled out a 2, used the quotient rule and now stuck factoring further

OpenStudy (anonymous):

what do I do with -4x/(x^2+1)^2

OpenStudy (bahrom7893):

i dont think u can do anything else with it..

OpenStudy (bahrom7893):

why?

OpenStudy (anonymous):

wolfram manages to break it down further

OpenStudy (anonymous):

polpak! I owe you answer still!

OpenStudy (bahrom7893):

are you sure the quotient rule was correct?

OpenStudy (anonymous):

Hrm?

OpenStudy (anonymous):

yes, has to be quotient rule

OpenStudy (anonymous):

\[\frac{d}{dx}[{(2-2x^2)\over (2+2x^2)}]\]\[ = \frac{d}{dx}[{(2-2x^2)(2+2x^2)^{-1}}]\]\[=(2-2x^2)[-1(2+2x^2)^{-2}(4x)] + [-4x](2+2x^2)^{-1}\]\[= {-4x(2-2x^2 + (2+2x^2)) \over (2+2x)^2}\]\[= {-16x \over (2+2x)^2} \]\[= {-16x \over [2(1+x^2)]^2}\]\[={-16x \over 4(x^2 +1)^2} = {-4x \over (x^2 +1)^2}\]

OpenStudy (anonymous):

that is exactly what i got

OpenStudy (anonymous):

my issue is that the program wolfram breaks it down further

OpenStudy (anonymous):

No it doesn't.

OpenStudy (bahrom7893):

yea.. it can't have..

OpenStudy (anonymous):

i had a capital X in the denom! haha! thanks for your help polpak - you rock out!